# UniFlow: Towards Zero-Shot LiDAR Scene Flow for Autonomous Vehicles via Cross-Domain Generalization

### v1

@@ 4.1 Ablation Studies
| + Augmentation | 15.76 ↓68.40 | 0.283 ↓0.048 | 0.039 ↑0.003 | 0.094 ↓0.025 | 0.127 ↓0.141 | 0.429 ↓0.230 |
Analysis on Dataset Weighting Strategies. We further examine the impact of per-dataset sampling frequencies on our unified training approach. All models are trained on the unified dataset for a fixed number of iterations, but we vary the per-domain sampling frequency to control how often samples from each distribution are drawn. As shown in Table 10, unified training consistently outperforms the single-dataset baselines, even under extreme weighting schemes. For example, on nuScenes, models trained with only 5% exposure to the target domain and 90% exposure to a very different distribution (Waymo) still surpass the in-domain baseline by 13.5% on Dynamic Mean EPE, indicating that even limited cross-domain exposure provides considerable benefits. Unsurprisingly, one can maximize in-domain performance by slightly biasing the model twoards in-domain data. Lastly, we find that uniformly sampling all datasets with equal probability yields the best out-of-distribution performance on TruckScenes.
@@
Analysis on Scaling Laws. We evaluate Flow4D (UniFlow) with varying amounts of training data on both in-distribution benchmarks (AV2, nuScenes, and Waymo) and an out-of-distribution benchmark (TruckScenes) in Fig 4. As expected, larger training sets reduce Dynamic Mean EPE. However, data augmentation proves far more critical for out-of-distribution performance, while its effect on in-distribution performance remains minimal.
| Method | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Dyn. Mean | Car | Other | Pedestr. | VRU | |
| 2 Hz | |||||||||
| Flow4D | 30.79 | 88.10 | 3.26 | 1.02 | 0.495 | 0.364 | 0.494 | 0.613 | 0.509 |
| SSF | 31.16 | 89.82 | 3.07 | 0.59 | 0.520 | 0.363 | 0.498 | 0.684 | 0.535 |
| Flow | 26.41 | 74.43 | 3.12 | 1.66 | 0.457 | 0.370 | 0.471 | 0.545 | 0.440 |
| Flow4D (UniFlow, Ours) | 27.52 | 79.54 | 2.36 | 0.65 | 0.444 | 0.297 | 0.415 | 0.587 | 0.476 |
| 5 Hz | |||||||||
| Flow4D | 9.58 | 25.64 | 2.45 | 0.65 | 0.323 | 0.215 | 0.312 | 0.432 | 0.333 |
| SSF | 8.74 | 23.02 | 2.74 | 0.47 | 0.311 | 0.179 | 0.286 | 0.489 | 0.291 |
| Flow | 6.43 | 16.54 | 1.98 | 0.76 | 0.264 | 0.173 | 0.271 | 0.369 | 0.242 |
| Flow4D (UniFlow, Ours) | 8.02 | 21.78 | 1.92 | 0.37 | 0.290 | 0.190 | 0.265 | 0.405 | 0.302 |
| 10 Hz (Standard Frame Rate) | |||||||||
| Flow4D | 3.46 | 7.97 | 1.85 | 0.55 | 0.230 | 0.160 | 0.241 | 0.241 | 0.176 |
| SSF | 3.00 | 6.55 | 2.04 | 0.41 | 0.220 | 0.142 | 0.197 | 0.398 | 0.144 |
| Flow | 2.05 | 4.19 | 1.36 | 0.59 | 0.173 | 0.132 | 0.184 | 0.261 | 0.117 |
| Flow4D (UniFlow, Ours) | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 |
| 20 Hz | |||||||||
| Flow4D | 2.55 | 5.79 | 1.35 | 0.50 | 0.272 | 0.196 | 0.315 | 0.377 | 0.200 |
| SSF | 2.94 | 6.72 | 1.73 | 0.38 | 0.316 | 0.177 | 0.303 | 0.533 | 0.250 |
| Flow | 1.60 | 3.28 | 0.48 | 0.48 | 0.220 | 0.167 | 0.231 | 0.327 | 0.156 |
| Flow4D (UniFlow, Ours) | 1.95 | 4.65 | 1.00 | 0.21 | 0.230 | 0.171 | 0.269 | 0.315 | 0.163 |
| Flow (UniFlow, Ours) | 1.30 | 2.60 | 0.96 | 0.34 | 0.010 | 0.153 | 0.215 | 0.277 | 0.138 |
Analysis on Frame-Rate. We evaluate the generalization of Flow4D (UniFlow), SSF (UniFlow) and Flow (UniFlow) across different frame rates in Table 11. Slower frame rates approximate a slower-moving ego-vehicle, while faster frame rates approximate a faster one. Although all models are trained using 10 Hz annotations, multi-dataset training consistently improves performance, especially at slower frame rates. Interestingly, higher frame rates have a smaller negative effect on performance, likely because models only need to estimate smaller displacement vectors for each timestep.
Limitations and Future Work. We investigate the domain gap between different LiDAR scene flow datasets and show that multi-dataset training enables strong generalization. Due to a lack of benchmarks and baselines in diverse domains, our experiments are limited to autonomous driving scenes. Future work should explore non-AV settings.
5 Conclusion
In this paper, we introduce UniFlow, a frustratingly simple approach that re-trains off-the-shelf LiDAR scene flow models with diverse data from multiple datasets. Although prior work in LiDAR-based 3D object detection and segmentation don’t seem to benefit from multi-dataset training, we posit that our proposed method works well “low-level” tasks such as motion estimation may be less sensitive to sensor configuration. Notably, our model establishes a new state-of-the-art Waymo and nuScenes, improving over prior work by 5.1% and 35.2% respectively. Although UniFlowis primarily trained on urban city driving scenarios, we demonstrate that it generalizes surprisingly well to highway truck driving without any dataset-specific fine-tuning. Our extensive analysis shows that such improvements are model-agnostic, suggesting that future scene flow methods should adopt a multi-dataset training strategy.
Appendix A Implementation Details
Models. We train and evaluate SSF and Flow using their official implementations without modification. As shown in Table 9, Flow4D benefits from a larger backbone when trained on the unified dataset. We denote this model as Flow4D-XL, which preserves the published architecture and increases all channel widths in the 4D Voxel Network by a uniform factor of three, with the Voxel Feature Encoders and Point Head adjusted accordingly to match the widened input/output dimensions. The full configuration is shown in Table 12. All dataset-specific Flow4D models use the original backbone, while all Flow4D models trained on the unified dataset use Flow4D-XL unless otherwise noted.
| Stage | STDB-P (Filters) | Pool & Up (Kernel, Stride) | Output shape |
| Input | – | – | |
| 1 | (48, 96) | Pool (2,2,2,1) | |
| (96, 96) | |||
| 2 | (96, 192) | Pool (2,2,2,1) | |
| (192, 192) | |||
| 3 | (192, 192) | Pool (2,2,2,1) | |
| (192, 192) | |||
| 4 | (192, 192) | Pool (2,2,1,1) | |
| (192, 192) | |||
| 5 | (192, 192) | Up (2,2,1,1) | |
| (192, 192) | |||
| 6 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 7 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 8 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 9 | (96, 48) | – |
@@ Appendix A Implementation Details
Augmentations. We apply two dataset-agnostic augmentations. Following Flow, we apply height jitter with probability 0.8 by translating the point cloud along the sensor’s z-axis using a uniform offset in the range of [0.5, 2.0] m. To improve cross-dataset density generalization, we also apply LiDAR beam dropout with probability 0.35, removing every other LiDAR beam for all Waymo and Argoverse 2 sweeps so that the effective density approximates a 32-beam sensor.
Training Settings. For the Argoverse 2 benchmark, we follow the standard evaluation protocol and report test-set performance according to the official public leaderboard to ensure fair comparison. For all TruckScenes experiments, baseline methods and their UniFlowvariants are trained under the same setups described in their respective original papers. UniFlow models adopt identical architectures unless otherwise noted, and all models are trained until convergence. We apply height augmentation to UniFlow and all dataset-specific Flow models to match the original training setup, and apply sparsity augmentation only to UniFlow models.
Appendix B Ablation on Synthetic Sparse LiDARs
We evaluate the performance of UniFlow models on synthetically sparsified LiDAR point clouds to investigate the impact of LiDAR sparsity on cross-domain generalization. We do not train any new models; all results are from models trained on the original datasets. Sparse datasets are created by removing every other beam from each LiDAR sweep. As shown in Table 13, the colored deltas indicate the change in performance relative to each model’s performance on the standard datasets (higher EPE indicates worse performance). Across datasets, degradation correlates strongly with dataset sparsity: models evaluated on Sparse nuScenes perform the worst due to its low (16-beam) density, while models evaluated on Sparse Waymo are affected the least due to its substantially higher point density.
Notably, UniFlow models are consistently more robust than dataset-specific baselines, exhibiting smaller performance drops across all datasets. This suggests that multi-dataset training provides useful regularization even under extreme sparsity.
@@ Appendix B Ablation on Synthetic Sparse LiDARs
| Flow4D (UniFlow, Ours) | 3.44 ↑0.20 | 8.27 ↑0.99 | 1.39 ↑0.13 | 0.65 ↑0.02 | 0.181 ↑0.005 | 0.138 ↑0.005 | 0.151 ↑0.002 | 0.198 ↓0.005 | 0.237 ↑0.017 |
@@
| Flow4D (UniFlow, Ours) | 1.66 ↑0.05 | 3.92 ↑0.09 | 0.73 ↑0.05 | 0.31 ↓0.02 | 0.143 ↓0.046 | 0.063 ↓0.001 | - | 0.249 ↓0.151 | 0.119 ↑0.015 |
@@
| Flow4D (UniFlow, Ours) | 3.72 ↑0.71 | 9.31 ↑2.03 | 1.52 ↑0.05 | 0.31 ↑0.03 | 0.279 ↑0.083 | 0.140 ↑0.003 | 0.266 ↑0.047 | 0.430 ↑0.158 | 0.279 ↑0.122 |
@@
Appendix C Ablation on Velocity Augmentation
@@
| Flow4D (UniFlow, Ours) | 0.191 | 0.064 | - | 0.400 | 0.110 | 0.006 | 0.053 | 0.042 | 0.128 |
| Flow4D (Fast-UniFlow, Ours) | 0.208 | 0.066 | - | 0.409 | 0.149 | 0.007 | 0.058 | 0.043 | 0.093 |
@@
| Flow4D (UniFlow, Ours) | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 | 0.009 | 0.082 | 0.089 | 0.374 |
| Flow4D (Fast-UniFlow, Ours) | 0.187 | 0.126 | 0.217 | 0.263 | 0.143 | 0.009 | 0.090 | 0.101 | 0.306 |
@@
| Flow4D (UniFlow, Ours) ZS | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 | 0.034 | 0.097 | 0.119 | 0.389 |
| Flow4D (Fast-UniFlow, Ours)ZS | 0.281 | 0.079 | 0.264 | 0.546 | 0.235 | 0.034 | 0.112 | 0.106 | 0.323 |
Appendix D Performance Breakdown on TruckScenes

### v2

@@ 4 Experiments
| + Augmentation | 16.04 ↓68.40 | 0.283 ↓0.048 | 0.039 ↑0.003 | 0.094 ↓0.025 | 0.127 ↓0.141 | 0.429 ↓0.230 |
| Method | Three-way EPE (cm) | Dynamic Bucket-Normalized | |||||||
| Mean | FD | FS | BS | Dyn. Mean | Car | Other | Pedestr. | VRU | |
| 2 Hz | |||||||||
| Flow4D | 30.79 | 88.10 | 3.26 | 1.02 | 0.495 | 0.364 | 0.494 | 0.613 | 0.509 |
| SSF | 31.16 | 89.82 | 3.07 | 0.59 | 0.520 | 0.363 | 0.498 | 0.684 | 0.535 |
| Flow | 26.41 | 74.43 | 3.12 | 1.66 | 0.457 | 0.370 | 0.471 | 0.545 | 0.440 |
| Flow4D-XL (UniFlow, Ours) | 27.52 | 79.54 | 2.36 | 0.65 | 0.444 | 0.297 | 0.415 | 0.587 | 0.476 |
| 5 Hz | |||||||||
| Flow4D | 9.58 | 25.64 | 2.45 | 0.65 | 0.323 | 0.215 | 0.312 | 0.432 | 0.333 |
| SSF | 8.74 | 23.02 | 2.74 | 0.47 | 0.311 | 0.179 | 0.286 | 0.489 | 0.291 |
| Flow | 6.43 | 16.54 | 1.98 | 0.76 | 0.264 | 0.173 | 0.271 | 0.369 | 0.242 |
| Flow4D-XL (UniFlow, Ours) | 8.02 | 21.78 | 1.92 | 0.37 | 0.290 | 0.190 | 0.265 | 0.405 | 0.302 |
| 10 Hz (Standard Frame Rate) | |||||||||
| Flow4D | 3.46 | 7.97 | 1.85 | 0.55 | 0.230 | 0.160 | 0.241 | 0.241 | 0.176 |
| SSF | 3.00 | 6.55 | 2.04 | 0.41 | 0.220 | 0.142 | 0.197 | 0.398 | 0.144 |
| Flow | 2.05 | 4.19 | 1.36 | 0.59 | 0.173 | 0.132 | 0.184 | 0.261 | 0.117 |
| Flow4D-XL (UniFlow, Ours) | 3.01 | 7.28 | 1.47 | 0.28 | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 |
| 20 Hz | |||||||||
| Flow4D | 2.55 | 5.79 | 1.35 | 0.50 | 0.272 | 0.196 | 0.315 | 0.377 | 0.200 |
| SSF | 2.94 | 6.72 | 1.73 | 0.38 | 0.316 | 0.177 | 0.303 | 0.533 | 0.250 |
| Flow | 1.60 | 3.28 | 0.48 | 0.48 | 0.220 | 0.167 | 0.231 | 0.327 | 0.156 |
| Flow4D-XL (UniFlow, Ours) | 1.95 | 4.65 | 1.00 | 0.21 | 0.230 | 0.171 | 0.269 | 0.315 | 0.163 |
| Flow (UniFlow, Ours) | 1.30 | 2.60 | 0.96 | 0.34 | 0.196 | 0.153 | 0.215 | 0.277 | 0.138 |
AEVAScenes Zero-Shot Performance. To further evaluate zero-shot generalization to unseen LiDARs, we benchmark our models on the AEVAScenes dataset 222https://scenes.aeva.com/, which features high-resolution FMCW LiDARs that differs substantially from rotating LiDAR sensors used in existing benchmarks. Due to the limited scale of AEVAScenes, all methods are evaluated zero-shot. Notably, UniFlow variants consistently outperform their respective best single-dataset counterparts on AEVAScenes, with SSF (UniFlow) improving Dynamic Mean EPE by 33.8%, Flow4D-XL (UniFlow) by 17.2%, and Flow (UniFlow) by 22.5%. We include visuals of zero-shot performance in Appendix 0.H.
Analysis of Long-Range Generalization. To evaluate whether unified training produces range-robust representations, we evaluate how models trained on short-range LiDAR data transfer to longer-range settings without retraining. Specifically, we extend the voxelization range during inference while keeping the voxel size identical to the original trained model. This change alters only the spatial coverage of the dynamic voxelization module, requiring a simple weight transfer step (i.e., model weights trained for short-range scene flow are loaded into a model with identical voxel size but a larger voxel grid for long-range processing) to adapt to the new grid dimensions while preserving all learned features. As shown in Table 8, FlowLR,ZS (UniFlow) consistently outperforms FlowLR across all distance intervals on AV2, reducing Dynamic Mean EPE by 12.7% on average and up to 20.0% in the 75–100 m range. This indicates that unified multi-dataset training produces representations that remain robust even when the perception range is significantly expanded. On TruckScenes, FlowLR,ZS (UniFlow) achieves a remarkable 40.6% reduction in overall error compared to FlowLR, with improvements exceeding 50% in the 50–100 m range, despite never being trained on this dataset. These results highlight that the representations learned through unified multi-dataset training naturally scale to larger spatial extents.
Analysis of Semantic Cross-Domain Generalization. We evaluate the benefits of unified training on both geometric and semantic understanding by extending Flow4D-XL with a multi-task semantic head. This new Flow4D-XL variant simultaneously predicts scene flow and per-point semantics using the same backbone features. Since all datasets follow different taxonomies, we map their annotations to four coarse categories (e.g. car, other, pedestrian, and VRU) and jointly train the model with an additional class-balanced cross-entropy loss. As shown in Table 9, adding the semantic head has almost no effect on scene flow accuracy, indicating that the auxiliary task does not interfere with motion estimation. Consistent with our original findings, unified multi-dataset training continues to improve both in-domain and zero-shot performance for scene flow. However, training the semantic prediction head on the unified dataset yields mixed results. Models trained on unified datasets show small in-domain gains, but perform substantially worse than dataset-specific model when evaluated on unseen domains. This supports our hypothesis that dataset unification primarily benefits low-level geometric representations, while high-level semantics remain sensitive to dataset-specific biases.
Analysis of Training Recipe. We ablate the impact of multi-dataset training, data augmentation, and model scaling on zero-shot TruckScenes performance in Table 10. All baselines are trained on Waymo only to ensure fair zero-shot evaluation. All model architectures show clear gains from multi-dataset training. Notably, despite training on a dataset three times larger, data augmentation continues to show substantial improvements, indicating that further scaling may yield additional performance gains.
Analysis on Scaling Laws. We evaluate Flow4D-XL (UniFlow) with varying amounts of training data on both in-distribution benchmarks (AV2, nuScenes, and Waymo) and an out-of-distribution benchmark (TruckScenes) in Fig 4. As expected, larger training sets reduce Dynamic Mean EPE. However, data augmentation proves far more critical for out-of-distribution performance, while its effect on in-distribution performance remains minimal.
Analysis on Frame-Rate. We evaluate the generalization of Flow4D-XL (UniFlow), SSF (UniFlow) and Flow (UniFlow) across different frame rates in Table 11. These settings serve as a stress test of temporal generalization outside standard operating conditions: slower frame rates approximate a slower-moving ego-vehicle, while faster frame rates approximate a faster one. Although all models are trained using 10 Hz annotations, multi-dataset training consistently improves performance over single-dataset models, especially at slower frame rates. Interestingly, higher frame rates have a smaller negative effect on performance, likely because models only need to estimate smaller displacement vectors for each timestep.
5 Conclusion
In this paper, we introduce UniFlow, a frustratingly simple approach for scalable LiDAR scene flow that re-trains off-the-shelf models with diverse data from multiple datasets. Although prior work in LiDAR-based 3D object detection and segmentation suggest that naively training on multiple LiDAR datasets yields worse results than single-dataset trained models, we posit that “low-level” tasks such as motion estimation may be less sensitive to sensor configuration. Notably, our model establishes a new state-of-the-art on Waymo and nuScenes, improving over prior work by 5.1% and 35.2% respectively. Although UniFlowis primarily trained on urban city driving scenarios and standard spinning LiDARs, we demonstrate that it generalizes surprisingly well to highway truck driving and novel FMCW sensors without any dataset-specific fine-tuning. Our extensive analysis shows that such improvements are model-agnostic, suggesting that future scene flow methods should adopt this multi-dataset training strategy.
Appendix 0.A Implementation Details
Models. We train and evaluate SSF and Flow using their official implementations without modification. As shown in Table 10, Flow4D benefits from a larger backbone when trained on the unified dataset. We denote this model as Flow4D-XL, which preserves the published architecture and increases all channel widths in the 4D Voxel Network by a uniform factor of three, with the Voxel Feature Encoders and Point Head adjusted accordingly to match the widened input/output dimensions. The full configuration is shown in Table 12. All dfataset-specific Flow4D models use the original backbone, while all Flow4D models trained on the unified dataset use Flow4D-XL unless otherwise noted. We train all models on 8x A6000 GPUs. Training takes 12 hours for baselines and 2 days for UniFlow. Runtime evaluations are conducted on a single A6000 GPU. Inference latency is identical for both approaches.
| Stage | STDB-P (Filters) | Pool & Up (Kernel, Stride) | Output shape |
| Input | – | – | |
| 1 | (48, 96) | Pool (2,2,2,1) | |
| (96, 96) | |||
| 2 | (96, 192) | Pool (2,2,2,1) | |
| (192, 192) | |||
| 3 | (192, 192) | Pool (2,2,2,1) | |
| (192, 192) | |||
| 4 | (192, 192) | Pool (2,2,1,1) | |
| (192, 192) | |||
| 5 | (192, 192) | Up (2,2,1,1) | |
| (192, 192) | |||
| 6 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 7 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 8 | (192, 192) | Up (2,2,2,1) | |
| (192, 192) | |||
| 9 | (96, 48) | – |
@@ Appendix 0.A Implementation Details
Datasets. We retain all scans from Waymo, Argoverse 2, nuScenes, and TruckScenes when constructing our unified dataset. For AEVAScenes, we manually filter the dataset to remove scenes with missing packets or unreliable ground-truth annotations, retaining 65 valid scenes out of 100. We use dataset provided ego motion for motion compensation.
@@
Augmentations. We apply two dataset-agnostic augmentations. Following Flow, we apply height jitter with probability 0.8 by translating the point cloud along the sensor’s z-axis using a uniform offset in the range of [-0.5, 2.0]m. To improve cross-dataset density generalization, we also apply LiDAR beam dropout with probability 0.35, removing every other LiDAR beam for all Waymo and Argoverse 2 sweeps to approximates a 32-beam sensor.
Training Settings. We follow the standard evaluation protocol and report test-set performance for the Argoverse 2 benchmark using the official public leaderboard for fair comparison. For all other datasets, baseline methods and their UniFlow variants are trained and evaluated on the prescribed train-val splits. Unless otherwise noted, UniFlow models use the same architectures as their corresponding baselines, and all models are trained until convergence. Flow and its UniFlow variants are trained on a m spatial grid with voxel resolution . Flow4D is trained on a m spatial range with voxel resolution , while SSF is trained on a m spatial range with voxel resolution . For the long-range generalization experiments in Table 8, we extend all evaluated models to a m range for each model while keeping each model’s voxel size fixed.
Appendix 0.B Analysis on Dataset Weighting Strategies.
We further examine the impact of per-dataset sampling frequencies on our unified training approach. All models are trained on the unified dataset for a fixed number of iterations, but we vary the per-domain sampling frequency to control how often samples from each distribution are drawn. As shown in Table 13, unified training consistently outperforms the single-dataset baselines, even under extreme weighting schemes. For example, on nuScenes, models trained with only 5% exposure to the target domain and 90% exposure to a very different distribution (Waymo) still surpass the in-domain baseline by 13.5% on Dynamic Mean EPE, indicating that even limited cross-domain exposure provides considerable benefits. Unsurprisingly, one can maximize in-domain performance by slightly biasing the model twoards in-domain data. Lastly, we find that uniformly sampling all datasets with equal probability yields the best out-of-distribution performance on TruckScenes.
@@
| Flow4D-XL (UniFlow, Ours) | 3.44 ↑0.20 | 8.27 ↑0.99 | 1.39 ↑0.13 | 0.65 ↑0.02 | 0.181 ↑0.005 | 0.138 ↑0.005 | 0.151 ↑0.002 | 0.198 ↓0.005 | 0.237 ↑0.017 |
@@
| Flow4D-XL (UniFlow, Ours) | 1.66 ↑0.05 | 3.92 ↑0.09 | 0.73 ↑0.05 | 0.31 ↓0.02 | 0.143 ↓0.046 | 0.063 ↓0.001 | - | 0.249 ↓0.151 | 0.119 ↑0.015 |
@@
| Flow4D-XL (UniFlow, Ours) | 3.72 ↑0.71 | 9.31 ↑2.03 | 1.52 ↑0.05 | 0.31 ↑0.03 | 0.279 ↑0.083 | 0.140 ↑0.003 | 0.266 ↑0.047 | 0.430 ↑0.158 | 0.279 ↑0.122 |
@@
Appendix 0.C Ablation on Synthetic Sparse LiDARs
We stress-test the performance of UniFlow models on synthetically sparsified LiDAR point clouds to investigate the impact of LiDAR sparsity on cross-domain generalization. We do not train any new models; all results are from models trained on the original datasets. Sparse datasets are created by removing every other beam from each LiDAR sweep. As shown in Table 14, the colored deltas indicate the change in performance relative to each model’s performance on the standard datasets (higher EPE indicates worse performance). Across datasets, degradation correlates strongly with dataset sparsity: models evaluated on sparse nuScenes perform the worst due to its low (16-beam) density, while models evaluated on sparse Waymo are affected the least due to its substantially higher point density. Notably, UniFlow models are consistently more robust than dataset-specific baselines, exhibiting smaller performance drops across all datasets. This suggests that multi-dataset training provides useful regularization even under extreme sparsity.
@@ Appendix 0.C Ablation on Synthetic Sparse LiDARs
| Flow4D-XL (UniFlow, Ours) | 0.191 | 0.064 | - | 0.400 | 0.110 | 0.006 | 0.053 | 0.042 | 0.128 |
| Flow4D-XL (Fast-UniFlow, Ours) | 0.208 | 0.066 | - | 0.409 | 0.149 | 0.007 | 0.058 | 0.043 | 0.093 |
@@
| Flow4D-XL (UniFlow, Ours) | 0.196 | 0.137 | 0.219 | 0.272 | 0.157 | 0.009 | 0.082 | 0.089 | 0.374 |
| Flow4D-XL (Fast-UniFlow, Ours) | 0.187 | 0.126 | 0.217 | 0.263 | 0.143 | 0.009 | 0.090 | 0.101 | 0.306 |
@@
| Flow4D-XL (UniFlow, Ours) ZS | 0.281 | 0.088 | 0.277 | 0.530 | 0.230 | 0.034 | 0.097 | 0.119 | 0.389 |
| Flow4D-XL (Fast-UniFlow, Ours)ZS | 0.281 | 0.079 | 0.264 | 0.546 | 0.235 | 0.034 | 0.112 | 0.106 | 0.323 |
Appendix 0.D Ablation on Velocity Augmentation